# AdS Length corrected Thermodynamics of a Black Hole in a Cloud of Strings and Perfect Fluid Dark Matter

### v1

@@ Abstract
The research paper investigates the thermodynamic behavior of a black hole in a cloud of strings (CoS) and perfect fluid dark matter (PFDM), incorporating a minor correction to the Anti-de Sitter (AdS) length scale. The study reveals that while the entropy remains unaffected by the AdS length correction, other thermodynamic quantities such as enthalpy, free energy, pressure, internal energy, Gibbs free energy, and heat capacity exhibit significant modifications. Positive corrections to the AdS length ($\epsilon>0$) generally enhance these quantities, leading to faster growth or slower decay with increasing black hole size, while negative corrections ($\epsilon<0$) often produce the opposite effect, sometimes resulting in divergent or negative values. These findings highlight the sensitivity of black hole thermodynamics to geometric corrections and exotic matter fields, providing deeper information about the interplay between such modifications and thermodynamic stability. These results underscore the importance of considering minor corrections in theoretical models to better understand black hole behavior in the presence of dark matter and string-like matter distributions.
Keywords: Cloud of Strings (CoS), Perfect Fluid Dark Matter (PFDM), AdS length correction, Quantum/ Thermal Corrections, Black Hole Thermodynamics.
1 Introduction
The study of black hole thermodynamics [1] in various modified gravity and matter field backgrounds has gained significant momentum in recent decades. This interest is driven both by the theoretical quest to unify general relativity with quantum field theory [2]-[10] and by astrophysical observations indicating the presence of dark matter and other exotic energy distributions in the universe.
Among the various matter models considered in the context of black hole spacetimes, the cloud of strings (CoS) and perfect fluid dark matter (PFDM) have emerged as compelling candidates. A cloud of strings represents a classical limit of string theory, offering a simplified model for string-like matter distributions that can arise in various high-energy scenarios [11]-[17]. On the other hand, PFDM has been proposed as a phenomenological model that effectively captures the gravitational influence of dark matter at both galactic and cosmological scales [18]-[24].
Incorporating such matter sources into black hole spacetimes leads to intriguing modifications in both geometric and thermodynamic properties. The presence of CoS and PFDM alters the black hole metric functions, influencing quantities like the horizon radius, surface gravity, and ultimately the Hawking temperature [25] and entropy [26].
Furthermore, within the framework of extended black hole thermodynamics, where the cosmological constant $\Lambda$ is treated as a thermodynamic pressure [27]-[31], the Anti-de Sitter (AdS) length scale $l$ plays a central role. Motivated by quantum corrections or higher-order effects, it is reasonable to consider minor modifications to the AdS length. These corrections, albeit small, can have significant implications for the thermodynamic behavior of the black hole, potentially affecting stability criteria, phase transitions, and critical phenomena [32]-[38]. Moreover, the minor corrections in AdS length have led to the universal relations [39].
In this paper, we investigate a static, spherically symmetric black hole solution surrounded by both a cloud of strings and perfect fluid dark matter [40]. We introduce a perturbative correction to the AdS length scale and study its impact on the thermodynamics of the system. Our goal is to analyze how such a correction influences the temperature, entropy, and heat capacity of the black hole, thereby shedding light on the interplay between the background matter fields and geometric corrections in AdS space.
The structure of the paper is as follows: In Section 2, we briefly review the black hole solution with CoS and PFDM in an AdS background. In Section 3, we incorporate a small correction to the AdS length and derive and analyze the modified thermodynamic quantities. We also analyze stability of the BH. We present our summarized results and outlook in Section 4.
@@ 2 Black Hole with Cloud of Strings and PFDM
The line element describing an AdS charged black hole (BH) in the presence of a cloud of strings (CoS) and perfect fluid dark matter (PFDM), as presented in [40], takes the following form:
@@
Here, $a$ represents the CoS parameter, $l$ is the AdS radius, $Q$ denotes the electric charge, and $\lambda$ characterizes the PFDM scale parameter.
Solving $f(r)=0$ at the event horizon radius $r_{+}$ yields the mass of the black hole:
$M_{+}=\frac{r_{+}}{2}\left(1+\frac{Q^{2}}{r_{+}^{2}}-a+\frac{r_{+}^{2}}{l^{2}}+\frac{\lambda}{2r_{+}}\ln\left(\frac{r_{+}}{\lambda}\right)\right)$ (1)
By evaluating the surface gravity at the horizon, the Hawking temperature is derived using $T_{H}=\frac{f^{\prime}(r_{+})}{4\pi}$:
@@
Incorporating the first law of thermodynamics for this system:
$dM_{+}=T_{+}dS_{+}+\Phi dQ+\lambda d\Lambda_{1}+PdV,$
where $\Lambda_{1}$ is conjugate to $\lambda$, and $\Phi$ is the electric potential.
@@
$S_{0}=\int\frac{1}{T_{+}}\frac{dM_{+}}{dr_{+}}dr_{+}=\pi r_{+}^{2}=\frac{A_{+}}{4}$ (3)
@@ 3 AdS length Corrected Thermodynamics
Figure 4: Internal energy $U_{c}$ v/s horizon radius $r_{+}$ with different values of $\epsilon$. The curves show the behavior for $\epsilon=0$ (uncorrected), $\epsilon>0$, and $\epsilon<0$.
@@
Figure 5: Gibbs free energy $G_{c}$ v/s horizon radius $r_{+}$ with different values of $\epsilon$. The curves show the behavior for $\epsilon=0$ (uncorrected), $\epsilon>0$, and $\epsilon<0$.
@@
Figure 6: Heat capacity $C_{c}$ v/s of horizon radius $r_{+}$ with different values of $\epsilon$. The curves show the behavior for $\epsilon=0$ (uncorrected), $\epsilon>0$, and $\epsilon<0$.
From the graph (6), we observe that the three curves converge for small horizon radii, being negligible, and depart from one another at $r_{+}\approx 0.3$. The curve corresponding to $\epsilon$ shows the same trend as the uncorrected curve but increases faster and becomes more positive with increasing black hole size and converges with it at large horizon radii. The negative correction exhibits divergence at $r_{+}\approx 0.75$ and gets closer to the other curves at large size.
@@ 4 Conclusions
The research paper investigated the thermodynamic behavior of a black hole in by a cloud of strings (CoS) and perfect fluid dark matter (PFDM), incorporating a minor correction to the Anti-de Sitter (AdS) length scale. The study revealed that while the entropy remains unaffected by the AdS length correction, other thermodynamic quantities such as enthalpy, free energy, pressure, internal energy, Gibbs free energy, and heat capacity exhibited significant modifications. Positive corrections to the AdS length ($\epsilon>0$) generally enhanced these quantities, leading to faster growth or slower decay with increasing black hole size, while negative corrections ($\epsilon<0$) often produced the opposite effect, sometimes resulting in divergent or negative values. These findings highlight the sensitivity of black hole thermodynamics to geometric corrections and exotic matter fields, providing deep information about the interplay between such modifications and thermodynamic stability. The results underscore the importance of considering minor corrections in theoretical models to better understand black hole behavior in the presence of dark matter and string-like matter distributions.

### v2

@@ Abstract
We investigate the thermodynamic characteristics of the Reisner-Nordstrom black hole in AdS spacetime with a cloud of strinsg and perfect fluid dark matter background, employing a minor correction to the AdS radius. Although the entropy remains unaltered, enthalpy, free energies, pressure, internal energy, and specific heat exhibit considerable modifications. These quantities get enhanced generally by the positive correction ($\epsilon>0$) to the AdS scale and lead to faster growth/ slower decay with increase in horizon radius. The opposite effect is reflected by negative corrections ($\epsilon<0$), which sometimes result in negative or divergent behaviour. We find an interconnection between the combined effect of exotic matter fields and geometric corrections on black hole thermodynamics and thermodynamic stability.
1 Introduction and Motivation
Black hole thermodynamics offers comprehensive understanding of gravity and spacetime, emerging as a crucial bridge between gravitation, statistical physics and quantum field theory [1, 2, 10]. Recent astrophysical observations have given boost to the field after signatures of dark matter and other exotic energy forms were found in the cosmos. Isolated black holes do not exist; rather they are often embedded in certain matter fields that affect their geometry and thermodynamics, in which the cloud of strings (CoS) and perfect fluid dark matter (PFDM) have been found to be strong candidates.
The CoS model represents a classical limit of string theory, describing string-like matter configurations that can naturally arise in high-energy scenarios [11, 17]. In contrast, the PFDM model was proposed to explain phenomena on cosmological scale such as galaxy rotation curves [18, 24]. Inclusion of such matter sources into the black hole solution affects fundamental properties of the likes of horizon radius, temperature[25] and entropy[26].
The interpretation of the cosmological constant $\Lambda$ as the thermodynamic pressure and its conjugate as the thermodynamic volume in the framework of black hole chemistry enriches the thermodynamic study of black holes [27, 31]. The AdS black holes undergo Hawking-Page and van der Waals-like phase transitions [32, 38]. The motivation for our current study to investigate minor corrections to the AdS radius as perturbation comes from the quantum gravity effects, higher order curvature corrections, renormalization of the cosmological constant etc., as these have demonstrated to modify critical conditions and discovery of universal thermodynamic relations [39, 42, 43], with persepective that minor geometric corrections can affect phase structures, stability and critical phenomena.
Even though black holes have been studied in detail in combined backgrounds of CoS and PFDM, and in various entropy correction schemes, much has not been worked upon AdS length corrected black holes in CoS and PFDM. This study will enhance our understanding of dependence of thermodynamic quantities like enthalpy, specific heat and free energies on the geometric perturbations.
In this manuscript, we study a spherically symmetric, static RN-AdS black embedded in a combined background of CoS and PFDM [44] by incorporating a correction to the AdS radius in the form of a perturbation. We analyze effects on the black hole’s entropy, temperature, specific heat and thermodynamic potentials.
We have organized the mansucript as follows. We present a brief overview of RN-AdS black holes embedded in a CoS and PFDM and evaluate the semi-classical thermodynamic paramters in sec. (2). We derive the corrected thermodynamic parameters, and anlyze thermodynamic stability, as a result of the introduction of the AdS length correction in sec. (3). And in sec. (4), we present summary of our results and future persepectives.
@@ 2 Black Hole with Cloud of Strings and PFDM
The solution for RN-AdS black hole embedded in a CoS and PFDM with respective parameters $a$ and $\lambda$, AdS length $l$ of the spacetime, charge $Q$ of the black hole is [44]
@@
The largest root of the equation $f(r_{+})=0$ at $r=r_{+}$ gives expression for the black hole mass
$M=\frac{r_{+}}{2}\left(1+\frac{Q^{2}}{r_{+}^{2}}-a+\frac{r_{+}^{2}}{l^{2}}+\frac{\lambda}{2r_{+}}\ln\left(\frac{r_{+}}{\lambda}\right)\right)$ (1)
Using $T_{H}=\frac{f^{\prime}(r_{+})}{4\pi}$, we get expression for the black hole temperature
@@
The first law of thermodynamics for the black hole in extended phase space is
$dM=T_{H}\;dS+\Phi\;dQ+\Pi_{\lambda}\;\lambda+\Xi_{a}\;da+V\;dP$ (3)
where $\Pi_{\lambda}$ denotes the PFDM potential, $\Xi_{a}$ is the CoS potential, and $\Phi$ is the electric potential.
@@
$S_{0}=\int\frac{1}{T_{H}}\frac{dM}{dr_{+}}dr_{+}=\pi r_{+}^{2}=\frac{A_{+}}{4}$ (4)
@@ 3 AdS length Corrected Thermodynamics
Figure 4: Corrected Internal energy $U_{c}$ v/s horizon radius $r_{+}$ with different values of $\epsilon$. The curves show the behavior for $\epsilon=0$ (uncorrected), $\epsilon>0$, and $\epsilon<0$.
@@
Figure 5: Corrected Gibbs free energy $G_{c}$ v/s horizon radius $r_{+}$ with different values of $\epsilon$. The curves show the behavior for $\epsilon=0$ (uncorrected), $\epsilon>0$, and $\epsilon<0$.
@@
Figure 6: Corrected Heat capacity $C_{c}$ v/s horizon radius $r_{+}$ with different values of $\epsilon$. The curves show the behavior for $\epsilon=0$ (uncorrected), $\epsilon>0$, and $\epsilon<0$.
From Fig. (6), we observe that the three curves converge for small horizon radii, the heat capacity values being negligible, implying instability of small-sized blacks holes, thus substantiating earlier results. The curves depart from one another at $r_{+}\approx 0.3$. The curve corresponding to $\epsilon$ shows the same trend as the uncorrected curve but increases faster and becomes more positive with increasing black hole size and converges with it at large horizon radii. The negative correction curve exhibits divergence in the heat capacity at the critical point $r_{+}\approx 0.75$, implying charged AdS black holes with CoS and PFDM undergo second-order Davies-type phase transition [45]. The curve gets closer to the other curves at large size, recovering classical behaviour.
@@ 4 Conclusions
In this research paper, we investigated the thermodynamic behavior of a black hole surrounded by a cloud of strings (CoS) and perfect fluid dark matter (PFDM), incorporating a minor correction to the Anti-de Sitter (AdS) length scale in the form of a perturbation. Though the correction did not affect entropy, but other thermodynamic quantities free energies, enthalpy, internal energy, specific heat, pressure, exhibited significant modification. These quantities were found to get enhanced by positive correction, resulting in faster growth/ slowed decay. On the other hand, negative corrections produce the contrasting effect, often giving negative/ divergent behaviour. Our results demonstrate the significance of minor AdS length corrections on thermodynamic behvaiour, stability and critical behaviour of Rn-AdS blacks in exotic matter fields.